OpenStudy (anonymous):

How do i use the substitution method to solve y=x+2 and 2x+y=8

OpenStudy (anonymous):

2x+(x+2)=8 3x=6 x=2

OpenStudy (lalaly):

substitute y=x+2 in 2x+y=8 2x+(x+2)=8 3x=6 x=2

OpenStudy (anonymous):

You have two equations y = x + 2 ->1 2x + y = 8 ->2 Now put the value of y in 1 in 2 2x + (x +2 ) = 8 3x =6 x =2 Then put x in 1 y = (2)+2 y=4

OpenStudy (anonymous):

thanks